# SWE-RL: Advancing LLM Reasoning via Reinforcement Learning on Open Software Evolution

### v1

@@ Abstract
The recent DeepSeek-R1 release has demonstrated the immense potential of reinforcement learning (RL) in enhancing the general reasoning capabilities of large language models (LLMs). While DeepSeek-R1 and other follow-up work primarily focus on applying RL to competitive coding and math problems, this paper introduces SWE-RL, the first approach to scale RL-based LLM reasoning for real-world software engineering. Leveraging a lightweight rule-based reward (e.g., the similarity score between ground-truth and LLM-generated solutions), SWE-RL enables LLMs to autonomously recover a developer’s reasoning processes and solutions by learning from extensive open-source software evolution data—the record of a software’s entire lifecycle, including its code snapshots, code changes, and events such as issues and pull requests. Trained on top of Llama 3, our resulting reasoning model, Llama3-SWE-RL-70B, achieves a 41.0% solve rate on SWE-bench Verified—a human-verified collection of real-world GitHub issues. To our knowledge, this is the best performance reported for medium-sized (<100B) LLMs to date, even comparable to leading proprietary LLMs like GPT-4o. Surprisingly, despite performing RL solely on software evolution data, Llama3-SWE-RL has even emerged with generalized reasoning skills. For example, it shows improved results on five out-of-domain tasks, namely, function coding, library use, code reasoning, mathematics, and general language understanding, whereas a supervised-finetuning baseline even leads to performance degradation on average. Overall, SWE-RL opens up a new direction to improve the reasoning capabilities of LLMs through reinforcement learning on massive software engineering data.
issue: ,issue: ,issue: ,
Figure 1: Overview of SWE-RL. We create a seed RL dataset from GitHub PR data, including issue descriptions, code context, and oracle patches. The policy LLM generates code changes through reasoning. For correctly formatted responses, the reward is calculated based on the match between the predicted and the oracle patch; incorrectly formatted responses are assigned a negative reward. GRPO is used for policy optimization.
@@ 1 Introduction
The application of large language models (LLMs) to software engineering (SE) tasks has received significant attention in recent years, with researchers exploring their potential to automate various complex SE tasks, such as library-level and complex code generation Zhuo et al. (2024); Zhao et al. (2024), real-world bug/issue resolution Xia and Zhang (2022); Jimenez et al. (2023); Yang et al. (2024c), and software testing Deng et al. (2023); Jain et al. (2024a). Among these tasks, SWE-bench Jimenez et al. (2023)—a benchmark for solving real-world software issues—has emerged as a focal point of research efforts, and researchers have proposed various agentic Yang et al. (2024b); Wang et al. (2024); Gauthier (2024) or pipeline-based Xia et al. (2024); Örwall (2024); Zhang et al. (2024) methods to push LLMs’ real-world issue solving capability. However, most current techniques depend on powerful proprietary LLMs like GPT-4o OpenAI (2024a) or Claude-3.5-Sonnet Anthropic (2024a), where advancements are driven more by enhanced prompting strategies than by improvements in the underlying LLM.
@@
With the release of DeepSeek-R1 DeepSeek-AI (2025), reinforcement learning (RL) using rule-based rewards has become a crucial technique for enhancing the reasoning capabilities of LLMs across various downstream tasks, including coding Zeng et al. (2025) and mathematics Yeo et al. (2025). However, their effectiveness in SE tasks remains limited DeepSeek-AI (2025), and their substantial total parameter size (671B) poses challenges for researchers attempting to train them. For mathematics, the reward is generally defined as whether the answer predicted by the LLM can exactly match the ground truth DeepSeek-AI (2025). While in coding, existing RL research typically utilizes execution feedback Gehring et al. (2025) as the reward signal and is limited to competitive programming tasks Li et al. (2022); Jain et al. (2024b), where code is self-contained and easily executable. This is challenging to apply to real-world SE tasks due to the execution cost and lack of executable environments Pan et al. (2024). Meanwhile, previous research Ma et al. (2024); Pan et al. (2024); Xie et al. (2025) relied on proprietary teacher models and has focused primarily on supervised fine-tuning (SFT), which, as we demonstrate in our paper, is less effective and less generalizable.
@@
To address these limitations, we propose SWE-RL, the first RL method to improve LLMs on SE tasks by directly using rule-based rewards and software evolution data—the record of entire software lifecycle, including all code snapshots, changes, and events like PRs and issues. As shown in Figure 1, we begin by curating a comprehensive dataset of GitHub pull requests (PRs), which is then transformed into the seed dataset for RL. Each data item includes an issue, the corresponding code context, and the oracle patch merged by the PR. During RL, the policy LLM is tasked with solving a given issue through reasoning and producing the code changes. The code changes are then converted into a consistent patch format for reward calculation. If the response is incorrectly formatted, the reward will be $-1$; otherwise, the reward is a similarity score (between 0 and 1) of the predicted and the oracle patch calculated by Python’s difflib.SequenceMatcher Ratcliff and Metzener (1988). Notably, we provide the complete content of each file in the input prompt, which implicitly teaches the model to reason about the precise fault locations before suggesting repair edits.
@@ 2 SWE-RL
2.1 Raw pull request data curation
Figure 2 provides a high-level overview of our process for curating the raw PR dataset for Llama3-SWE-RL. In the following paragraphs, we detail each step of the curation process. During data processing, we exclude all the repositories used by SWE-bench Jimenez et al. (2023) to prevent data contamination.
PR data aggregation. These collected events and git clones are disparate entities that require further processing before they can be used for training purposes. At this stage, we focus on each PR individually and aggregate all pertinent information associated with it. This includes mentioned issues, user discussions, review comments, initial code contents, and subsequent commits and code changes.
To start, we keep only merged PRs and gather all related conversational events for each PR, sorting them in chronological order. Next, using the base_commit and head_commit hashes of a PR, we retrieve the contents of all modified files indicated by its patch at the merge base of the two commits. The reasoning behind this approach is that many PRs aim to merge back into the main branch, which may have undergone changes since the PR was created. By considering the merge base as the actual starting point for a developer working on the PR, we can more accurately understand the context of the changes. We save all intermediate commits and code changes made between the merge base and the head commit. Additionally, we extract the complete patch that represents the cumulative changes from start to finish. Finally, we scan each aggregated PR to identify patterns that resemble issues, and associate the matched issues with the corresponding PR. In the end, we have 24M aggregated PR instances.
Relevant files prediction. Currently, each pull request includes only the code files that have been modified. In our earlier experiments, we noticed that this approach let LLMs learn a bias: the model consistently generated edits for every code file presented and was unable to handle noisy files presented in the context. This issue was also mentioned in the finetuning experiment discussed in the SWE-bench paper Jimenez et al. (2023). To mitigate this problem, we prompt Llama-3.1-70B-Instruct Dubey et al. (2024) to generate a list of relevant but unmodified files for each PR, given the PR description and the paths of the changed files, and include the contents of these files in our final dataset.
Data filtering. GitHub PRs can be quite noisy, so we implemented various filtering strategies to eliminate potentially harmful PRs. In designing these filtering rules, our goal is to maximize the recall of high-quality PRs while permitting a certain level of noise. First, we remove the bot-generated PRs whose title, description, or username contains keywords “[bot]”, “dependabot”, “renovate”, “bump”, or “automerge”. Also, we remove PRs with empty changes or with extremely large number of changes (e.g., in some PRs, the developer uploaded a directory of data files by mistake). Additionally, we implemented a more fine-grained set of filters used in CodeLlama Rozière et al. (2023) to examine each code change hunk. We then removed any PRs where all code changes were flagged by these filters. For example, this can exclude PRs with only lock file changes or version updates. Finally, this gives us around 11M unique PR instances.
@@ 2.2 Reward modeling
Figure 3: Prompt template used to train Llama3-SWE-RL with SWE-RL. Given an issue description and the corresponding code context, the policy LLM needs to generate search/replace edits Xia et al. (2024) to fix this issue through reasoning. This is the only subtask we incorporate in the RL training. During inference, the LLM can generalize to tasks outside the training domain (e.g, file-level localization and test generation). For conciseness, we exclude certain prompt details, with the complete prompt template available in Section 9.
To prepare the initial dataset for RL, we extract high-quality PR seeds from the raw dataset we collected. These seeds are selected based on specific heuristics. For example, a PR instance should include at least one linked issue, the issue should describe a bug-fixing request, and the code changes should involve programming files. For each seed, we extract the issue descriptions and code context, including all changed files and some relevant but unchanged files. These are converted to input prompts for the policy LLM. We also take the oracle patch from each seed, which will be used for reward calculation.
@@
We bootstrap the policy LLM with the prompt template shown in Figure 3. Assuming that the LLM has generated a rollout $\tau$ given an issue and its code context, the reward function is defined as follows:
[…]
Here, $\mathsf{patch_{pred}}$ denotes the patch corresponding to the code changes generated by the policy LLM if the format of $\tau$ is correct, and $\mathsf{patch_{gt}}$ means the oracle patch for this issue.
[…]
@@
Given a seed RL dataset $\mathcal{D}_{\mathsf{seed}}{}$ where each item contains (1) , the issue description, (2) $\mathsf{ctx}$, the code context required to solve this issue, including both files to repair and relevant files not intended for editing, and (3) $\mathsf{patch_{gt}}{}$, which is the oracle patch. The input prompt for each data item is formed by instantiating the prompt template in Figure 3 with the issue description and code context, which we denote as $q=\mathsf{form\mbox{-}prompt}(\mathsf{ctx})$. The policy LLM $\mathit{\pi_{\theta}}$ tries to solve the issue by generating code changes through reasoning, and produces multiple outputs $o_{i}$ for each input prompt $q$ given the group size $G$. Then, the policy LLM aims to maximize the following GRPO objective:
@@
$\begin{gathered}\mathcal{J}(\theta)={\mathbb{E}}\left[\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(A/_{i},\mathrm{clip}\left(,/1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta D_{\mathrm{KL}}(\mathit{\pi_{\theta}}\,\|\,\mathit{\pi_{\mathrm{ref}}})\right)\right],\\ \text{where }(\mathsf{ctx},\mathsf{patch_{gt}})\sim\mathcal{D}_{\mathsf{seed}}\text{, }q=\mathsf{form\mbox{-}prompt}(\mathsf{ctx})\text{, and }\{o_{i}\}_{i=1}^{G}\sim\mathit{\pi_{\theta_{\mathrm{old}}}}(\cdot\mid q).\end{gathered}$ (2)
@@
In this equation, $\epsilon$ and $\beta$ are hyperparameters, and $\mathit{\pi_{\theta_{\mathrm{old}}}}$ and $\mathit{\pi_{\mathrm{ref}}}$ are the old and reference policy, respectively. Following GRPO, the advantages $A_{i}$ are calculated using the normalized rewards within each group. The term $D_{\mathrm{KL}}$ denotes the approximated KL-divergence Schulman (2020).
@@
Our training approach conditions the model on the complete context of each file, implicitly forcing it to identify detailed fault locations before generating repair edits. This process inherently teaches the model both bug diagnosis and repair generation. However, during downstream evaluation on SWE-bench, our Agentless Mini scaffold requires capabilities beyond generating repair edits, such as file-level fault localization, reproduction test generation, and regression test selection. Remarkably, even without explicit training on these subtasks, Llama3-SWE-RL can generalize to them through the RL process.
@@ 2.3 Aha moments and generalized reasoning capabilities
“Aha moments” on software engineering. With the application of SWE-RL, we observe “aha moments” DeepSeek-AI (2025) where Llama3-SWE-RL exhibits emergent reasoning skills. To our knowledge, this is the first study demonstrating the existence of such aha moments in the realm of real-world software engineering tasks, confirming the findings of DeepSeek-R1 DeepSeek-AI (2025), which mainly focuses on competition coding and math. As shown in Figure 4, through RL, Llama3-SWE-RL-70B can allocate more thinking time to reflect on its initial assumptions during the issue-solving process. This behavior emerges naturally from the model’s interaction with RL, rather than being explicitly programmed.
General reasoning capabilities. Surprisingly, we have identified additional aha moments where Llama3-SWE-RL acquires general reasoning abilities that are transferrable to various out-of-domain tasks, such as function-level code generation and mathematics, although RL is applied exclusively to software issue solving. Figure 4 demonstrates that Llama3-SWE-RL is capable of reasoning through self-reflection, exploring alternative approaches, and solving complex problems by breaking them down into smaller subtasks. In Section 3.5, we demonstrate that Llama3-SWE-RL improves over even more out-of-domain tasks, including library use, code reasoning, and general language understanding.
@@ 3.1 Experimental setup
Training configs. Llama3-SWE-RL-70B is trained on top of Llama-3.3-70B-Instruct Dubey et al. (2024) using SWE-RL for 1,600 steps with a 16k context window. We use a global batch size of 512, sampling 16 rollouts from each of the 32 problems in every batch. For every global step, a single optimization step is performed.
Scaffolding. We have developed Agentless Mini on top of Agentless Xia et al. (2024) as the underlying scaffold. Different from Agentless’s multi-step localization, Agentless Mini focuses solely on file-level localization, delegating detailed reasoning to the repair step by providing the entire file contents in the input. This enables Llama3-SWE-RL to do more reasoning during SWE-RL and simplifies the RL process to focus on only one issue-solving task. Despite this simplification, Llama3-SWE-RL can still seamlessly complete other pipeline steps, benefiting from out-of-domain generalization through RL. Furthermore, while Agentless only employs one reproduction test per issue for reranking, Agentless Mini can use multiple reproduction tests, which has proven effective in our scaling analysis (Section 3.4). More details about Agentless Mini can be found in Section 6.
Evaluation setup. We conduct evaluation on SWE-bench Verified OpenAI (2024), a subset of SWE-bench with 500 human-verified problems that can more reliably evaluate AI models’ capability in solving real-world software issues. In the main evaluation (Section 3.2), we generate 500 patches for each problem using a 1.0 temperature, and use the top 30 reproduction tests for execution and reranking. Ultimately, only the highest-ranked patch for each issue will be submitted for SWE-bench evaluation to calculate the pass@1 score.
SFT baseline. To understand the advantages of SWE-RL, we also trained an SFT baseline, named Llama3-SWE-SFT-70B, for experiments in Section 3.3, Section 3.4, and Section 3.5. It is trained on top of Llama-3.3-70B-Instruct Dubey et al. (2024) using a mixture of synthetic code editing data, Llama 3 Dubey et al. (2024) coding SFT data, and Llama 3 general SFT data. The synthetic data is generated using an approach inspired by Magicoder Wei et al. (2024), where high-quality PR data serves as the seeds for creating chain-of-thoughts and subsequent editing, as well as serving as the oracle for filtering. Llama-3.3-70B-Instruct is employed for this generation process. The seed PR dataset is identical to the one utilized for RL. In contrast to our RL model, which only requires a seed PR dataset to trigger the RL loop, the SFT baseline needs synthetic data generation for chain-of-thoughts and additional data mix to ensure the dataset diversity and model generalizability. More details are explained in Section 7.

### v2

@@ Abstract
The recent DeepSeek-R1 release has demonstrated the immense potential of reinforcement learning (RL) in enhancing the general reasoning capabilities of large language models (LLMs). While DeepSeek-R1 and other follow-up work primarily focus on applying RL to competitive coding and math problems, this paper introduces SWE-RL, the first approach to scale RL-based LLM reasoning for real-world software engineering. Leveraging a lightweight rule-based reward (e.g., the similarity score between ground-truth and LLM-generated solutions), SWE-RL enables LLMs to autonomously recover a developer’s reasoning processes and solutions by learning from extensive open-source software evolution data—the record of entire software development cycles, including code snapshots, code changes, and events such as issues and pull requests. Trained on top of Llama 3, our resulting reasoning model, Llama3-SWE-RL-70B, achieves a 41.0% solve rate on SWE-bench Verified—a human-verified collection of real-world GitHub issues. To our knowledge, this is the best performance reported for medium-sized (<100B) LLMs to date, even comparable to leading proprietary LLMs like GPT-4o. Surprisingly, despite performing RL solely on software evolution data, Llama3-SWE-RL has even emerged with generalized reasoning skills. For example, it shows improved results on five out-of-domain tasks, namely, function coding, library use, code reasoning, mathematics, and general language understanding, whereas a supervised-finetuning baseline even leads to performance degradation on average. Overall, SWE-RL opens up a new direction to improve the reasoning capabilities of LLMs through reinforcement learning on massive software engineering data.
Figure 1: Overview of SWE-RL. We create a seed RL dataset from GitHub PRs, including issue descriptions, code context, and oracle patches. A policy LLM generates code edits via reasoning. Rewards are based on the similarity to oracle patches, with penalties for formatting errors. We optimize the policy using GRPO.
@@ 1 Introduction
The application of large language models (LLMs) to software engineering (SE) tasks has received significant attention, with researchers exploring their potential to automate various complex SE tasks, such as library-level and complex code generation Zhuo et al. (2024); Zhao et al. (2024), real-world bug/issue resolution Xia and Zhang (2022); Jimenez et al. (2023); Yang et al. (2024c), and software testing Deng et al. (2023); Jain et al. (2024a). Among these tasks, SWE-bench Jimenez et al. (2023)—a benchmark for solving real-world software issues—has emerged as a focal point of research efforts, and researchers have proposed various agentic Yang et al. (2024b); Wang et al. (2024b); Gauthier (2024) and pipeline-based Xia et al. (2024); Örwall (2024); Zhang et al. (2024) methods to push LLMs’ real-world issue solving capability. However, most current techniques rely on powerful proprietary LLMs like GPT-4o OpenAI (2024a) or Claude-3.5-Sonnet Anthropic (2024a), where advancements are driven more by enhanced prompting strategies than by improvements in the underlying LLM.
@@
With the release of DeepSeek-R1 DeepSeek-AI (2025), reinforcement learning (RL) using rule-based rewards has become a crucial technique for enhancing the reasoning capabilities of LLMs across various downstream tasks, including coding Zeng et al. (2025) and mathematics Yeo et al. (2025). However, their effectiveness in SE tasks remains limited DeepSeek-AI (2025), and their substantial total parameter size (671B) poses challenges for researchers attempting to train them. For mathematics, the reward is generally defined as whether the answer predicted by the LLM can exactly match the ground truth DeepSeek-AI (2025). While in coding, existing RL research typically utilizes execution feedback DeepSeek-AI (2025); Gehring et al. (2025) as the reward signal and is limited to competitive programming tasks Li et al. (2022); Jain et al. (2024b), where code is self-contained and easily executable. This is challenging to apply to real-world SE tasks due to the execution cost and lack of executable environments Pan et al. (2024). Meanwhile, previous research Ma et al. (2024); Pan et al. (2024); Xie et al. (2025) relied on proprietary teacher models and focused primarily on supervised fine-tuning (SFT), which, as we show in the paper, is less effective and less generalizable.
@@
To address these limitations, we propose SWE-RL, the first RL method to improve LLMs on SE tasks by directly using rule-based rewards and software evolution data—the record of entire software lifecycle, including all code snapshots, changes, and events like PRs and issues. As shown in Figure˜1, we begin by curating a comprehensive dataset of GitHub pull requests (PRs), which is then transformed into the seed dataset for RL. Each data item includes an issue, the corresponding code context, and the oracle patch merged by the PR. During RL, the policy LLM is tasked with solving a given issue through reasoning and producing the code changes. The code changes are then converted into a consistent patch format for reward calculation. If the response is incorrectly formatted, the reward will be 1-1; otherwise, the reward is a similarity score (between 0 and 1) of the predicted and the oracle patch calculated by Python’s difflib.SequenceMatcher Ratcliff and Metzener (1988). Notably, we provide the complete content of each file in the input prompt, which implicitly teaches the model to reason about the precise fault locations before suggesting repair edits.
@@ 2 SWE-RL
To prepare the initial dataset for RL, we extract 273k high-quality PR seeds from the raw PR dataset we collected from GitHub. These seeds are selected based on specific heuristics. For example, a PR instance should include at least one linked issue, the issue should describe a bug-fixing request, and the code changes should involve programming files. For each seed, we extract the issue descriptions and code context, including all changed files and some relevant but unchanged files. These are converted to input prompts for the policy LLM. We also take the oracle patch from each seed, which will be used for reward calculation. More details are explained in Appendix˜A.
@@ 2.1 Reward modeling
Figure 2: Prompt template used to train Llama3-SWE-RL with SWE-RL. Given an issue description and the corresponding code context, the policy LLM needs to generate search/replace edits Xia et al. (2024) to fix this issue through reasoning. This is the only subtask we incorporate in the RL training. During inference, the LLM can generalize to tasks outside the training domain (e.g, file-level localization and test generation). For conciseness, we exclude certain prompt details, with the complete prompt template available in Appendix˜D.
@@
We bootstrap the policy LLM with the prompt template shown in Figure˜2. Assuming that the LLM has generated a rollout $o$ given an issue and its code context, the reward function is defined as follows:
$\mathcal{R}(o)=\begin{cases}-1,&\text{if}\ o\ \text{has wrong format},\\ \mathit{compare}(\mathsf{patch_{pred}},\mathsf{patch_{gt}}),&\text{otherwise}.\end{cases}$ (1)
Here, $\mathsf{patch_{pred}}$ is the patch extracted from the LLM generation $o$ if it is correctly formatted, and $\mathsf{patch_{gt}}$ means the oracle patch for this issue.
In the implementation, we use Python’s difflib.SequenceMatcher as the $\mathit{compare}$ function, which returns a floating point between 0 and 1 indicating the sequence similarity, and adopt Group Relative Policy Optimization (GRPO) Shao et al. (2024) for policy optimization.
@@
Given a seed RL dataset $\mathcal{D}_{\mathsf{seed}}{}$ where each item contains (1) $\mathsf{issue}{}$, the issue description, (2) $\mathsf{ctx}$, the code context required to solve this issue, including both files to repair and relevant files not intended for editing, and (3) $\mathsf{patch_{gt}}{}$, which is the oracle patch. The input prompt for each data item is formed by instantiating the prompt template in Figure˜2 with the issue description and code context, which we denote as $q=\mathsf{form\mbox{-}prompt}(\mathsf{issue},\mathsf{ctx})$. The policy LLM $\mathit{\pi_{\theta}}$ tries to solve the issue by generating code changes through reasoning, and produces multiple outputs $o_{i}$ for each input prompt $q$ given the group size $G$. Then, the policy LLM aims to maximize the following GRPO objective:
@@
$\small\begin{gathered}\mathcal{J}(\theta)={\mathbb{E}}\left[\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(\frac{\mathit{\pi_{\theta}}(o_{i}\mid q)}{\mathit{\pi_{\theta_{\mathrm{old}}}}(o_{i}\mid q)}A_{i},\mathrm{clip}\left(\frac{\mathit{\pi_{\theta}}(o_{i}\mid q)}{\mathit{\pi_{\theta_{\mathrm{old}}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta D_{\mathrm{KL}}(\mathit{\pi_{\theta}}\,\|\,\mathit{\pi_{\mathrm{ref}}})\right)\right],\\ \text{where }(\mathsf{issue},\mathsf{ctx},\mathsf{patch_{gt}})\sim\mathcal{D}_{\mathsf{seed}}\text{, }q=\mathsf{form\mbox{-}prompt}(\mathsf{issue},\mathsf{ctx})\text{, and }\{o_{i}\}_{i=1}^{G}\sim\mathit{\pi_{\theta_{\mathrm{old}}}}(\cdot\mid q).\end{gathered}$ (2)
@@
$\epsilon$ and $\beta$ are hyperparameters, and $\mathit{\pi_{\theta_{\mathrm{old}}}}$ and $\mathit{\pi_{\mathrm{ref}}}$ are the old and reference policy. The rewards $r_{i}=\mathcal{R}(o_{i})$ and advantages $A_{i}=\frac{r_{i}-\operatorname{mean}(r_{1},\ldots,r_{G})}{\operatorname{std}(r_{1},\ldots,r_{G})}$ are calculated using the normalized rewards within each group following GRPO. $D_{\mathrm{KL}}$ denotes the estimated KL-divergence Schulman (2020).
@@
Our training approach conditions the model on the complete context of each file, implicitly forcing it to identify detailed fault locations before generating repair edits. This process inherently teaches the model both bug diagnosis and repair generation. However, during the evaluation on SWE-bench, Agentless Mini requires capabilities beyond generating repair edits, such as file-level fault localization, reproduction test generation, and regression test selection. Remarkably, even without explicit training on these subtasks, Llama3-SWE-RL can generalize to them through the RL process.
@@ 2.2 Aha moments and generalized reasoning capabilities
“Aha moments” on software engineering. With the application of SWE-RL, we observe “aha moments” DeepSeek-AI (2025) where Llama3-SWE-RL exhibits emergent reasoning skills. To our knowledge, this is the first study demonstrating the existence of such aha moments in the realm of real-world software engineering tasks, confirming the findings of DeepSeek-R1 DeepSeek-AI (2025), which mainly focuses on competition coding and math. As shown in Figure˜3, through RL, Llama3-SWE-RL-70B can allocate more thinking time to reflect on its initial assumptions during the issue-solving process. This behavior emerges naturally from the model’s interaction with RL, rather than being explicitly programmed.
General reasoning capabilities. Surprisingly, we have identified additional aha moments where Llama3-SWE-RL acquires general reasoning abilities that are transferrable to various out-of-domain tasks, such as function-level code generation and mathematics, although RL is applied exclusively to software issue solving. Figure˜3 demonstrates that Llama3-SWE-RL is capable of reasoning through self-reflection, exploring alternative approaches, and solving complex problems by breaking them down into smaller subtasks. In Section˜3.5, we demonstrate that Llama3-SWE-RL improves over even more out-of-domain tasks, including library use, code reasoning, and general language understanding.
@@ 3.1 Experimental setup
Training configs. Llama3-SWE-RL-70B is trained on top of Llama-3.3-70B-Instruct Dubey et al. (2024) using SWE-RL for 1,600 steps with a 16k context window. We use a global batch size of 512, sampling 16 rollouts from each of the 32 problems in every batch. For every global step, a single optimization step is performed using Adam Kingma and Ba (2017). We train our models on 512 NVIDIA H100 GPUs; a training run takes approximately 32 wall-time hours.
Scaffolding. We have developed Agentless Mini on top of Agentless Xia et al. (2024) as the underlying scaffold. Different from Agentless’s multi-step localization, Agentless Mini focuses solely on file-level localization, delegating detailed reasoning to the repair step by providing the entire file contents in the input. This enables Llama3-SWE-RL to do more reasoning during SWE-RL and simplifies the RL process to focus on only one issue-solving task. Despite this simplification, Llama3-SWE-RL can still seamlessly complete other pipeline steps, benefiting from out-of-domain generalization through RL. Furthermore, while Agentless only employs one reproduction test per issue for reranking, Agentless Mini can use multiple reproduction tests, which has proven effective in our scaling analysis (Section˜3.4). More details about Agentless Mini can be found in Appendix˜B.
Evaluation setup. We conduct evaluation on SWE-bench Verified OpenAI (2024), a subset of SWE-bench with 500 human-verified problems that can more reliably evaluate AI models’ capability in solving real-world software issues. In the main evaluation (Section˜3.2), we generate 500 patches for each problem using a 1.0 temperature, and use the top 30 reproduction tests for execution and reranking. Ultimately, only the highest-ranked patch for each issue will be submitted for SWE-bench evaluation to calculate the pass@1 score.
SFT baseline. To understand the advantages of SWE-RL, we also trained an SFT baseline, named Llama3-SWE-SFT-70B, for experiments in Section˜3.3, Section˜3.4, and Section˜3.5. It is trained on top of Llama-3.3-70B-Instruct Dubey et al. (2024) using a mixture of synthetic code editing data, Llama 3 Dubey et al. (2024) coding SFT data, and Llama 3 general SFT data. The synthetic data is generated using an approach inspired by Magicoder Wei et al. (2024), where high-quality PR data serves as the seeds for creating chain-of-thoughts and subsequent editing, as well as serving as the oracle for filtering. Llama-3.3-70B-Instruct is employed for this generation process. In contrast to our RL model, which only requires a seed PR dataset to trigger the RL loop, the SFT baseline needs synthetic data generation for chain-of-thoughts and additional data mix to ensure the dataset diversity and model generalizability. More details are explained in Appendix˜C.